OpenStudy (anonymous):

log_4(x+4)-log_16x=1 solve

sam (.sam.):

\[\log _{4}(x+4)-\log _{16}x=1\] \[\log _{4}(x+4)-\frac{\log _{4}x}{\log _{4}4^{2}}=1\] \[\log _{4}(x+4)-\frac{\log _{4}x}{2}=1\] \[\log _{4}(x+4)^{2}-\log _{4}x=2\] \[\log _{4}\frac{(x+4)^{2}}{x}=2\] \[\frac{(x+4)^2}{x}=4^2\] \[x^2+8x+16=16x\] \[x^2+8x=0\] \[x=0~,~x=-8\]

sam (.sam.):

x^2-8x+16=0 x=4